A market has a demand curve of Pd= 25 - 4Q and a supply curve of PS= 15 + Q

A. Illustrate graphically.

B. Calculate the consumer and producer surplus.

C. Whom is benefiting in this market?

D. Identity the firm type?
Answer #1

We have the following information

Demand equation: P = 25 – 4Q

Supply equation: P = 15 + Q

In the above P is price and Q is quantity

We will equate the demand and supply equations to get the equilibrium quantity

25 – 4Q = 15 + Q

5Q = 10

Equilibrium quantity = 2

P = 15 + Q

P = 15 + 2

Equilibrium price = 17

For the demand equation

When Q = 0, P = 25 and

When P = 0, Q = 6.25

For the supply equation

When Q = 0, P = 15 and

When P = 0, Q = – 15

Consumer surplus is the difference between the highest price a consumer is willing to pay and the actual price paid by the consumer. Consumer surplus is given by the area of the triangle ABC

Area of the triangle ABC = ½ × Base × Height

Base = 25 – 17 = 8

Height = 2

Consumer surplus = ½ × 8 × 2

Consumer surplus = 8

Producer surplus is the difference between the actual price of a good or service, which the market price, and the lowest price a producer is willing to accept for the good. Producer surplus is given by the area of the triangle ABD

Area of the triangle ABD = ½ × Base × Height

Base = 17 – 15 = 2

Height = 2

Producer surplus = ½ × 2 × 2

Producer surplus = 2

The producer surplus is less than the consumer surplus. From this one can say that we are dealing with a perfectly competitive market.

brice 29 Consumu Surplus supply curve 17B Producer Susplus 15 < Demand Equation T 1 TAS o Q 6.25 Quantity